# Where Are All The Tourists From 3025?

### v1

@@ I Introduction
On 28 June 2009, Stephen Hawking held a party for time-travellers [1, 2, 3]. Champagne, balloons, nibbles: this party had all the trappings of an era-defining party, hoping to entice time-travellers into making an appearance. However, nobody was invited to the event – in advance – and it was only publicly announced afterwards, with exact time and location details, in the hope that an invite would survive the centuries/millennia required to fall into the hands of a party-loving time-traveller. Despite all the effort put into throwing and publicizing this party, “no one came” [4]. Leaving Hawking to party alone.
@@ II Preliminary Analysis and Model Setup
The model I use to analyze time-travel as a phenomena is derived from a full description of the many possible worlds [10] it is possible to be living in (i.e. the set of worlds that could have been the world if things had been different). The worlds are each modeled by a full logical description of them via a – potentially infinite – set of logical statements that are true in that world. The worlds are then grouped into sets of worlds based on those that are the same in the regard I care about, the number of time machines that are ever constructed, and this gives rise to the true model I analyse herein.
@@ II.3 Dynamics of the Model: Symmetries and State Transitions
Additionally, I assume transitions are entirely symmetric but only between macrostates with construction numbers that differ by one. I.e., the transitions probabilities are only non-zero between macrostates where the difference between their construction numbers is exactly one and, in this case, the probability of a transition only depends on the macrostate the transition is from. I.e. if the construction number is increasing or decreasing does not affect the transition rate.
@@ II.4 A Formal Statistical Model
To formally define my model, I begin by constructing a Markov chain, where each state of the Markov chain is a macrostate (as defined in Def. 4444I.e. the state space of the Markov chain is the set of all Macrostates and the transition probabilities are proportional to the transition rates referred to above (scaling by the duration each time step lasts in the Markov chain). The aforementioned time step duration is then allowed to tend to zero, to obtain a continuous time Markov chain and lose the unphysical discretizations.
@@
Therefore, the probability of a macrostate transitioning to itself (i.e. not transitioning) is:
$1-\beta\dfrac{j}{2(j+1)}-\beta\dfrac{j}{2(j+1)}$ $=\dfrac{1+(1-\beta)j}{j+1}.$ (5)
@@ III.2 Perceptions of the Effects of Time Travel for Non-Travellers
So, factoring in the discussion of orthogonal time in Sec. III, the only timeline we should ever, as non-travellers, expect to see is the asymptotic limit of orthogonal time. That is, in the upcoming analysis or when I run my simulations the most important feature will be what happens in the limit as orthogonal time tends to infinity. As the only time that means anything when considering the dynamics of timelines (and hence the one I have to use in my model) is orthogonal time. So the limit as the travellers orthogonal time tends to infinity is all non-travellers (i.e. us) will ever see. All the intermediate states of the timeline – from changes that were undone or themselves changed – are erased (or more accurately, are prevented from ever having happened) and only the asymptotic limit timeline remains.
@@ IV.1 Reformulation as a System of Differential Equations
I now aim to find the dynamics of the model constructed in the previous section. The most important quantities of the model are the probabilities of being in each state; to aid in the analysis of these quantities, I define some notation for them in the below Def. 9.
@@
$\dfrac{dP_{j}^{t}}{dt}$ $=\alpha_{j-1}P_{j-1}^{t}+\alpha_{j+1}P_{j+1}^{t}-\alpha_{j}P_{j}^{t},$ (6)
@@
$\Delta P_{j}^{t}$ $=\bigg{[}\alpha_{j-1}P_{j-1}^{t-\Delta t}+\alpha_{j+1}P_{j+1}^{t-\Delta t}-\alpha_{j}P_{j}^{t-\Delta t}\bigg{]}\Delta t$ (7)
@@
$\lim_{\Delta t\longrightarrow 0}\bigg{(}\dfrac{\Delta P_{j}^{t}}{\Delta t}\bigg{)}$ $=\lim_{\Delta t\longrightarrow 0}\bigg{(}\alpha_{j-1}P_{j-1}^{t-\Delta t}+\alpha_{j+1}P_{j+1}^{t-\Delta t}-\alpha_{j}P_{j}^{t-\Delta t}\bigg{)}$ (8)
$\Rightarrow\dfrac{dP_{j}^{t}}{dt}$ $=\alpha_{j-1}P_{j-1}^{t}+\alpha_{j+1}P_{j+1}^{t}-\alpha_{j}P_{j}^{t}.$ (9)
@@
I note that the system of differential equations derived in Lemma 6 is as large as the model is, and so for large models is large. It is useful to note that all equations in the system are linear and that there is the additional requirement that: $\forall t\in\mathbb{R}^{+}$,
@@ IV.2 Asymptotic Behavior of the Model
Consider two times $t^{\prime}\in\mathbb{R}^{+}$ and $t\in\mathbb{R}^{+}$, such that $t^{\prime}>t$. I can express the probability that the model is in state $j\in\mathbb{N}_{0}$ at time $t^{\prime}$ by the probability it is in state $j$ at time $t$ plus the probability it moves into state $j$ in the intervening time (between time $t^{\prime}$ and time $t$) minus the probability it moves into state $j$ in the intervening time. When $j=0$, this is slightly simpler as the probability of moving out of state 0 at any time is zero and therefore $\mathbb{P}\big{(}X_{t^{\prime}}=0\big{)}$ (i.e. $P^{t^{\prime}}_{0}$) can be expressed as: $\forall t^{\prime}>t\in\mathbb{R}^{+}$,
@@
I then, in a slight diversion, reason that (as per Def. 8):
$\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$ $,\alpha_{j}\not=0$ (12)
$\Rightarrow\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$ $,\mathbb{P}\big{(}X_{t^{\prime}}=0\text{ }|\text{ }X_{t}=j\big{)}>0.$ (13)
Therefore, returning to Eqn. 11 and using Eqn. 13, in addition to the fact that $\mathbb{P}\big{(}X_{t^{\prime}}=0\text{ }|\text{ }X_{t}=0\big{)}=1$,
@@ V.2 System with 500 Macrostates, Starting in Macrostate 3
V.3 System with 500 Macrostates, Starting in Macrostate 250
@@ VI Comparison to Other Suggested Explanations
The most obvious solution to us not seeing travellers from the future is also the least interesting: time travel may not be possible for technical reasons (i.e. physics does not permit travel backwards in time). It may simply not be possible to travel through time. This may be the true solution – Occam’s razor would suggest so – but it is not definitively the case so other explanations are explored below and compared to the model originated herein.

### v2

@@ I Introduction
On 28 June 2009, Stephen Hawking held a party for time-travellers [1, 2]. Champagne, balloons, nibbles: this party had all the trappings of an era-defining party, hoping to entice time-travellers into making an appearance. However, nobody was invited to the event – in advance – and it was only publicly announced afterwards, with exact time and location details, in the hope that an invite would survive the centuries/millennia required to fall into the hands of a party-loving time-traveller. Despite all the effort put into throwing and publicizing this party, “no one came” [3]. Leaving Hawking to party alone.
@@ II Preliminary Analysis and Model Setup
The model I use to analyze time-travel as a phenomenon is derived from a full description of the many possible worlds [9] it is possible to be living in (i.e. the set of worlds that could have been the world if things had been different). The worlds are each modeled by a full logical description of them via a – potentially infinite – set of logical statements that are true in that world. The worlds are then grouped into sets of worlds based on those that are the same in the regard I care about, the number of time machines that are ever constructed, and this gives rise to the true model I analyse herein.
@@ II.3 Dynamics of the Model: Symmetries and State Transitions
Additionally, I assume that, for any fixed macrostate, the only possible transitions are to macrostates with construction numbers that differ from the construction number of the fixed macrostate by one. The transition rates for the possible transitions out of the fixed state are all identical. I.e., the transitions probabilities are only non-zero between macrostates where the difference between their construction numbers is exactly one and, in this case, the probability of a transition only depends on the macrostate the transition is from [^3] Meaning if the construction number is increasing or decreasing does not affect the transition rate.
@@ II.4 A Formal Statistical Model
To formally define my model, I begin by constructing a Markov chain, where each state of the Markov chain is a macrostate (as defined in Def. 4555I.e. the state space of the Markov chain is the set of all Macrostates and the transition probabilities are proportional to the transition rates referred to above (scaling by the duration each time step lasts in the Markov chain). The aforementioned time step duration is then allowed to tend to zero, to obtain a continuous time Markov chain and lose the unphysical discretizations.
@@ III.2 Perceptions of the Effects of Time Travel for Non-Travellers
So, factoring in the discussion of orthogonal time in Sec. III, the only timeline we should ever, as non-travellers, expect to see is the asymptotic limit of orthogonal time. As the model takes place entirely in orthogonal time, in the upcoming analysis or when I run my simulations the most important feature will be what happens in the limit as orthogonal time tends to infinity. As the only time that means anything when considering the dynamics of timelines (and hence the one I have to use in my model) is orthogonal time. So the limit as the travellers orthogonal time tends to infinity is all non-travellers (i.e. us) will ever see. All the intermediate states of the timeline – from changes that were undone or themselves changed – are erased (or more accurately, are prevented from ever having happened) and only the asymptotic limit timeline remains.
Note 1.
I pause to note that the “time” in the Markov chains in the model I will construct is itself orthogonal time, as just defined.
@@ IV.1 Reformulation as a System of Differential Equations
I now aim to find the dynamics of the model constructed in the previous section. These dynamics take place as orthogonal time passes. The most important quantities of the model are the probabilities of being in each state; to aid in the analysis of these quantities, I define some notation for them in the below Def. 9.
@@
$\dfrac{dP_{j}^{t}}{dt}$ $=\alpha_{j-1}P_{j-1}^{t}+\alpha_{j+1}P_{j+1}^{t}-2\alpha_{j}P_{j}^{t},$ (5)
@@
$\Delta P_{j}^{t}$ $=\bigg[\alpha_{j-1}P_{j-1}^{t-\Delta t}+\alpha_{j+1}P_{j+1}^{t-\Delta t}-2\alpha_{j}P_{j}^{t-\Delta t}\bigg]\Delta t$ (6)
@@
$\lim_{\Delta t\longrightarrow 0}\bigg(\dfrac{\Delta P_{j}^{t}}{\Delta t}\bigg)$ $=\lim_{\Delta t\longrightarrow 0}\bigg(\alpha_{j-1}P_{j-1}^{t-\Delta t}+\alpha_{j+1}P_{j+1}^{t-\Delta t}-2\alpha_{j}P_{j}^{t-\Delta t}\bigg)$ (7)
$\Rightarrow\dfrac{dP_{j}^{t}}{dt}$ $=\alpha_{j-1}P_{j-1}^{t}+\alpha_{j+1}P_{j+1}^{t}-2\alpha_{j}P_{j}^{t}.$ (8)
@@
I note that for Eqn. 5 to hold for all values of $j\geq 0$ (i.e. all states in the Markov chain) requires:
  1. 1.
    I additionally – purely so the variables in Eqn. 5 are always defined – define, $\forall t\in\mathbb{R}^{+}$, $\alpha_{-1}=P_{-1}^{t}=0$ and $\alpha_{\text{Max}+1}=P_{\text{Max}+1}^{t}=0$.
  2. 2.
    Implicitly, $\alpha_{\text{Max}}$ is the only transition rate that depends on the direction of the transition (i.e. if the construction number is increasing or decreasing), if the transition is to state $\text{Max}+1$, $\alpha_{\text{Max}}=0$. Otherwise, it is as in Def. 8.
  3. 3.
    The system of differential equations derived in Lemma 1 (and given as Eqn. 5) is as large as the model is, and so for large models is large. It is useful to note that all equations in the system are linear and that there is the additional requirement that: $\forall t\in\mathbb{R}^{+}$,
@@ IV.2 Asymptotic Behavior of the Model
Consider any two times $t^{\prime}\in\mathbb{R}^{+}$ and $t\in\mathbb{R}^{+}$, such that $t^{\prime}>t$.
I can express the probability that the model is in state $j\in\mathbb{N}_{0}$ at time $t^{\prime}$ as:
the probability the model is already in state $j$ at time $t$
plus
the probability it moves into state $j$, from any other state, after time $t$ but before time $t^{\prime}$
minus
the probability it moves out of state $j$ after time $t$ but before time $t^{\prime}$.
When $j=0$, this analysis is slightly simpler as the probability of moving out of state 0 at any time is zero and therefore $\mathbb{P}\big(X_{t^{\prime}}=0\big)$ (i.e. $P^{t^{\prime}}_{0}$) can be expressed as: $\forall t^{\prime}>t\in\mathbb{R}^{+}$,
@@
In a slight diversion, consider that, as $\forall j\in\mathbb{W}^{\prime}\backslash\{0\},\alpha_{j}>0$$\forall t^{\prime},t\in\mathbb{R}$ such that $t^{\prime}-t>0$,
$\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$ $,\mathbb{P}\big(X_{t^{\prime}}=0\text{ }|\text{ }X_{t}=j\big)>0.$ (11)
This implies that, as, $\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$, $\mathbb{P}\big(X_{t}=j\big)\geq 0$ and $\sum_{j=0}^{\text{Max}}\big(\mathbb{P}\big(X_{t}=j\big)\big)=1$: $\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$,
$\sum_{j\in\mathbb{W}^{\prime}\backslash\{0\}}\bigg(\mathbb{P}\big(X_{t^{\prime}}=0\text{ }|\text{ }X_{t}=j\big)\mathbb{P}\big(X_{t}=j\big)\bigg)>0,$ (12)
unless $\forall j\in\mathbb{W}^{\prime}\backslash\{0\}$, $\mathbb{P}\big(X_{t}=j\big)=0$.
Therefore, returning to Eqn. 10 and using Eqn. 12, in addition to the fact that $\mathbb{P}\big(X_{t^{\prime}}=0\text{ }|\text{ }X_{t}=0\big)=1$,
@@ V.2 System with 500 Macrostates, Starting in Macrostate 3
V.3 System with 500 Macrostates, Starting in Macrostate 100
@@ VI Comparison to Other Suggested Explanations
The most obvious solution to us not seeing travellers from the future is also the least interesting: time travel may not be possible for technical reasons (i.e. physics does not permit travel backwards in time [^8] Although results already exist to indicate reversing the flow of time is possible [29, 30].). It may simply not be possible to travel through time. This may be the true solution – Occam’s razor would suggest so – but it is not definitively the case so other explanations are explored below and compared to the model originated herein.
@@
Note 2.
After the first version this paper had been uploaded to arXiv, I was made aware that Larry Niven had suggested an informal version of the conclusion of this paper in Ref. [26], which he terms Niven’s law:
“In the universe of discourse permits the possibility of time travel and of changing the past, then no time machine will be invented in that universe.”
@@ VII Conclusions and Discussion
VIII Acknowledgements
I would like to thank James Lyons-Weiler, for extensive comments on the first version of this paper, and James Smirle, for making me aware of Niven’s Law.